# Membership Inference Attack with Partial Features

### v1

@@ IV-B Detailed Attack Procedure
Memory-Guided Feature Reconstruction: As outlined in SectionIV-A, we reconstruct the unknown features by minimizing the loss of the model on the sample. As shown in Algorithm 1, for a sample with missing feature values, we first initialize unknown features (line 1). Since the reconstruction process relies on information from the known features, it is important that the model maintains its reliance on them after initialization, in order to prevent the initial values of the unknown features from dominating the early gradient updates. To this end, we initialize all unknown features to zero. We then combine the known features and the initialized unknown features according to their original order(line 2). Since the attacker have access to the loss function during the training process, then we can iteratively update the unknown features via backpropagation to minimize the loss of the sample (lines 3–7).
Anomaly Detection: As shown in Algorithm 2, to assess the degree of deviation of a sample, we first model the distribution of real data (lines 1–8). This involves sampling an auxiliary inference set $\mathcal{D}_{aux}$ from the data distribution and computing the centroid $c_{k}$ and dispersion for each class. For measuring dispersion, we adopt an approach similar to CADE [22], using the Median Absolute Deviation (MAD) as a robust indicator of spread. Then we compute the distance $\hat{d}$ between the reconstructed sample $\hat{x}_{k}$ and the centroid of class $k$, and derive the corresponding base deviation distance $\delta$ (lines 9–10). To further enhance inference performance, we train a shadow model $\theta_{s}$ on data drawn from the same distribution and compute a shadow deviation distance $\delta_{s}$ on shadow model based on the same known features, following the same reconstruction procedure. In summary, both $\delta$ and $\delta_{s}$ are derived using identical feature values, but from the target and shadow models respectively. The shadow deviation $\delta_{s}$ serves as a reference for how typical the deviation would be if the known features is not a member of the training set. Then, we define a membership inference rule based on the relative deviation ${\delta_{s}}/{\delta}$. Let MRAD denote this relative deviation and the inference rule can be formulated as:
@@
Threshold Selection: Previous work has shown that MIA is only meaningful when the false positive rate is sufficiently low[5]. Therefore, we adopt a threshold selection strategy similar to that proposed by Ye et al.[14]. Specifically, the adversary can sample a batch of samples $\mathcal{D}_{non}$ from the overall non-member data space $\pi^{\prime}$. Given a predefined acceptable false positive rate $\alpha$, the adversary then selects a threshold that satisfies
@@
Algorithm 1 Memory-guided Reconstruction
0: Known features $x_{kno}$ from class k, target model $\theta$, number of iterations $N_{itr}$, step length $\eta$, known feature mask $M$, loss function $\mathcal{L}$ during the training process of the target model.
[…]
3:for $i=1$ to $N_{itr}$ do
4:  $g\leftarrow grad(\mathcal{L}_{\theta}(\hat{x},k))$
5:  $(g_{kno},g_{unk})\leftarrow\text{split}(g,M)$ // Split the gradient vector into gradients of known features and gradients of unknown features.
6:  $x_{unk}\leftarrow x_{unk}-\eta\cdot g_{unk}$
7:  $\hat{x}\leftarrow\text{cat}(x_{kno},x_{unk},M)$
8:end for
9:return  $\hat{x}$
@@ V Evaluation
Next, we evaluate the attack performance of our proposed algorithm. We begin by presenting the experimental setup, followed by a comprehensive assessment of the attack performance with widely used benchmark datasets. We then analyze the sensitivity of algorithm to hyperparameters and examine how the importance of known features influence performance. Finally, we conduct a case study to demonstrate the practical applications of our attack.
@@ V-A Experiment setup
Dataset & Model: We evaluate our attack with three widely used image datasets and one tabular dataset. For all image datasets, the target model is AlexNet, while for the tabular dataset, we employ a multilayer perceptron (MLP). We train each model for 500 epochs with a learning rate of 0.001. We provide a detailed overview of each dataset as below.
  • CIFAR-10[37]: CIFAR-10 is a widely used benchmark dataset for image classification. It contains 50,000 training images and 10,000 test images, each being a 32×32 RGB image.
  • Fashion-MNIST[38]: Fashion-MNIST is a drop-in replacement for the original MNIST dataset, featuring 60,000 training images and 10,000 test images. Each image is a 28×28 grayscale representation of an article of clothing.
  • STL-10[39]: STL-10 contains 13,000 labeled images and 100,000 unlabeled images, making it suitable for supervised, unsupervised, and self-supervised learning. In our experiments, we use only the labeled subset. Each image is a 96×96 RGB image.
  • Epsilon[40]: Epsilon was derived from the PASCAL 2008 challenge, contains simulated physical data for a binary classification task. Each sample consists of 2,000 normalized numerical features.
@@
Anomaly Detection Methods: In addition to our proposed method, we integrate three representative anomaly detection techniques: CADE [22], MSAD [25], and NCI [26]. These comparisons demonstrate the compatibility of our framework with various existing techniques and allow for a comprehensive evaluation across different methodologies. The three baseline methods are summarized as follows:
@@
  • NCI: NCI is a post-hoc OOD detection method based on the phenomenon of neural collapse. It observes that in-distribution (ID) samples tend to align with their class prototype vectors in the penultimate layer, whereas OOD samples do not. Additionally, ID features tend to cluster near the origin. We adopt the publicly available implementation from github[41].
V-B Intuitive Experiment
To offer a preliminary assessment of the effectiveness of our method, we first perform a simple validation experiment based on our core design intuition. In this experiment, we compute the basic deviation $\delta$ using the target model $\theta$ for both member and non-member features and compare the results. For each dataset, we randomly designate half of the features as known. Because the gradient magnitude is small, we set a larger step size of $\eta=100$ and perform only one optimization step (the effect of hyperparameters will be further discussed in Section V-D). As shown in Figure3, across all four datasets, member samples consistently show much lower deviation values than non-member samples. These results support the soundness of our intuition and provide initial evidence for the viability of the proposed attack.
(a) CIFAR-10
(b) STL-10
(c) Fashion-MNIST
(d) Epsilon
Figure 3: Distribution of the base deviation $\delta$ for member (blue) and non-member (red) samples across different datasets.
V-C Performance Evaluation
We begin by evaluating the attack performance under different proportions of known features. We vary the known feature percentage from 10% to 90% in increments of 10% and conduct experiments at each level. In addition, we include a leave-one-out setting where only a single feature is unknown. The results are shown in Figures 4 and 5, where the x-axis indicates the percentage of known features. The leave-one-out setting is marked as ”loo” on the x-axis. We also plot the full ROC curves on a log scale presented in Figure 6. The evaluation results indicate that the attack begins to demonstrate meaningful effectiveness when the known feature percentage reaches 50%. As the proportion of known features increases, an overall improvement is observed in both evaluation metrics. Notably, even the simple anomaly detection method proposed in this work delivers strong performance. In particular, it consistently achieves higher true positive rates under low false positive rate conditions across most datasets.
We attribute the upward trend to the increased amount of information available to the attacker as the proportion of known features grows. With more known features, the optimization space becomes more constrained, allowing the attack easier to converge. However, we observe a notable performance drop when the known feature percentage reaches 90%. This decline is likely due to the limited number of unknown features. Since the known features remain unchanged, a high known percentage leads to reconstructed samples looks like real ones. In such cases, the known features dominate the anomaly detection process, making it harder for the model to capture meaningful deviations. At 100% known features, no modification is needed, which results in the failure of all attack methods. Consequently, the task reduces to a standard Membership Inference Attack, for which conventional MIA techniques can be applied directly.
@@ V-C Performance Evaluation
(a) CIFAR-10 60% KP
(b) CIFAR-10 70% KP
(c) CIFAR-10 80% KP
(d) CIFAR-10 90% KP
(e) STL-10 60% KP
(f) STL-10 70% KP
(g) STL-10 80% KP
(h) STL-10 90% KP
(i) Fashion-MNIST 60% KP
(j) Fashion-MNIST 70% KP
(k) Fashion-MNIST 80% KP
(l) Fashion-MNIST 90% KP
(m) Epsilon 60% KP
(n) Epsilon 70% KP
(o) Epsilon 80% KP
(p) Epsilon 90% KP
Figure 6: Log-scale ROC curves on different datasets under varying known feature percentages (KP: Known Percentage).
V-D Hyperparameter Sensitivity
To have a better understanding of the behavior of our attack, we further analyze the influence of two key hyperparameters with a known percentage of 80%.
Performance VS Number of Iterations: We first assess the sensitivity of our algorithm to the number of optimization iterations, $N_{\text{itr}}$. With the step size $\eta$ fixed at 100, we vary $N_{\text{itr}}$ and report the corresponding AUC and TPR@0.1FPR, as shown in Table I and Table II. The results indicate that the performance of our algorithm remains largely stable across different iteration counts, suggesting low sensitivity to this parameter. Remarkably, even a single iteration is sufficient to achieve strong performance.
TABLE I: AUC performance under different numbers of optimization iterations ($N_{itr}$) across various datasets.
$N_{itr}$ CIFAR-10 STL-10 Fashion-MNIST Epsilon
Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI
1 0.658 0.670 0.672 0.652 0.636 0.705 0.668 0.738 0.590 0.576 0.583 0.615 0.662 0.667 0.701 0.735
2 0.641 0.655 0.660 0.684 0.670 0.694 0.683 0.714 0.577 0.582 0.566 0.606 0.656 0.674 0.690 0.723
5 0.627 0.668 0.659 0.669 0.675 0.697 0.687 0.701 0.570 0.598 0.571 0.590 0.634 0.673 0.710 0.748
10 0.645 0.647 0.668 0.643 0.664 0.690 0.682 0.690 0.567 0.575 0.579 0.582 0.643 0.679 0.723 0.719
20 0.623 0.647 0.676 0.665 0.677 0.694 0.690 0.687 0.567 0.601 0.569 0.599 0.649 0.687 0.725 0.742
TABLE II: TPR@0.1FPR under different numbers of optimization iterations ($N_{itr}$) across various datasets.
$N_{itr}$ CIFAR-10 STL-10 Fashion-MNIST Epsilon
Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI
1 0.283 0.261 0.252 0.217 0.264 0.314 0.268 0.324 0.192 0.117 0.128 0.177 0.244 0.157 0.184 0.241
2 0.295 0.260 0.243 0.286 0.328 0.280 0.266 0.333 0.172 0.124 0.127 0.187 0.237 0.162 0.155 0.212
5 0.273 0.259 0.264 0.266 0.371 0.284 0.315 0.326 0.158 0.133 0.127 0.147 0.238 0.161 0.175 0.254
10 0.265 0.243 0.254 0.259 0.341 0.309 0.318 0.310 0.162 0.124 0.123 0.152 0.227 0.182 0.211 0.201
20 0.240 0.232 0.261 0.282 0.382 0.308 0.314 0.332 0.179 0.119 0.132 0.157 0.229 0.204 0.222 0.221
Performance vs. Step Length: Next, we evaluate how the step length parameter $\eta$ affects the attack performance. The results are summarized in Table III and Table IV. In this experiment, we fix the number of optimization iterations to $N_{\text{itr}}=1$. We find that when $\eta$ is too small, the attack becomes less effective. This is because samples with known features exhibit low gradient magnitudes. As $\eta$ increases, attack performance improves consistently across various datasets and settings.
TABLE III: AUC performance of our attack on different datasets under varying step lengths ($\eta$).
$\eta$ CIFAR-10 STL-10 Fashion-MNIST Epsilon
Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI
0.01 0.488 0.495 0.518 0.497 0.489 0.517 0.477 0.484 0.471 0.500 0.507 0.500 0.541 0.495 0.498 0.509
0.1 0.457 0.496 0.521 0.493 0.478 0.517 0.477 0.486 0.484 0.503 0.504 0.505 0.512 0.482 0.496 0.485
1 0.460 0.490 0.513 0.498 0.457 0.516 0.471 0.488 0.536 0.538 0.535 0.517 0.481 0.457 0.504 0.553
10 0.623 0.645 0.530 0.635 0.466 0.486 0.483 0.610 0.588 0.580 0.570 0.607 0.605 0.595 0.659 0.701
100 0.656 0.678 0.654 0.635 0.657 0.706 0.672 0.728 0.593 0.580 0.557 0.592 0.648 0.666 0.702 0.758
1000 0.639 0.667 0.680 0.670 0.663 0.713 0.716 0.723 0.606 0.595 0.578 0.588 0.680 0.711 0.763 0.777
TABLE IV: TPR@0.1FPR of our attack on different datasets under varying step lengths ($\eta$).
$\eta$ CIFAR-10 STL-10 Fashion-MNIST Epsilon
Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI Ours CADE MSAD NCI
0.01 0.097 0.108 0.103 0.093 0.096 0.116 0.077 0.095 0.074 0.089 0.108 0.111 0.155 0.090 0.081 0.110
0.1 0.062 0.089 0.105 0.096 0.075 0.102 0.084 0.098 0.079 0.097 0.093 0.126 0.092 0.079 0.095 0.091
1 0.074 0.105 0.097 0.110 0.075 0.101 0.071 0.101 0.092 0.111 0.092 0.123 0.026 0.077 0.092 0.117
10 0.260 0.207 0.112 0.210 0.084 0.080 0.098 0.189 0.189 0.115 0.119 0.172 0.146 0.111 0.160 0.189
100 0.302 0.222 0.255 0.227 0.279 0.313 0.274 0.296 0.147 0.119 0.119 0.157 0.233 0.153 0.166 0.236
1000 0.271 0.244 0.248 0.275 0.347 0.328 0.320 0.291 0.177 0.139 0.137 0.146 0.291 0.147 0.248 0.281
V-E Ablation Study

### v2

@@ IV-B Detailed Attack Procedure
Memory-Guided Feature Reconstruction: As shown in Algorithm 1, for each sample with missing feature values, we begin by randomly sampling a label $k$ from label set $K$ (line 1) to account for the fact that a single feature combination may be associated with multiple labels. We then initialize all unknown features to zero (line 2). This zero-initialization ensures that, during the early stages of optimization, the model’s predictions remain primarily influenced by the known features rather than arbitrary initialization noise. Next, we merge the known and initialized unknown features into their original ordering to form a complete input vector (line 3). Since the attacker has access to the loss function used during the model’s training, the unknown features can be iteratively optimized through backpropagation (lines 4–8). By repeatedly updating the unknown features to minimize the sample’s training loss, the procedure gradually reconstructs feature values that are consistent with the target model’s learned decision boundary.
Algorithm 1 Memory-guided Reconstruction
0: Known features $x_{kno}$ from class k, possible label index set $K$, target model $\theta$, number of iterations $N_{itr}$, step length $\eta$, known feature mask $M$, loss function $\mathcal{L}$ during the training process of the target model.
[…]
4:for $i=1$ to $N_{itr}$ do
5:  $g\leftarrow grad(\mathcal{L}_{\theta}(\hat{x},k))$
6:  $(g_{kno},g_{unk})\leftarrow\text{split}(g,M)$ // Split the gradient vector into gradients of known features and gradients of unknown features.
7:  $x_{unk}\leftarrow x_{unk}-\eta\cdot g_{unk}$
8:  $\hat{x}\leftarrow\text{cat}(x_{kno},x_{unk},M)$
9:end for
10:return $\hat{x}$
Anomaly Detection: As shown in Algorithm 2, to assess the degree of deviation of a sample, we first model the distribution of real data (lines 1–8). This involves sampling an auxiliary inference set $\mathcal{D}_{aux}$ from the data distribution and computing the centroid $c_{k}$ and dispersion for each class. For measuring dispersion, we adopt an approach similar to CADE [yang2021cade], using the Median Absolute Deviation (MAD) as a robust indicator of spread. Then we compute the distance $\hat{d}$ between the reconstructed sample $\hat{x}_{k}$ and the centroid of class $k$, and derive the corresponding base deviation distance $\delta$ (lines 9–10). To further enhance inference performance, we train a shadow model $\theta_{s}$ on data drawn from the same distribution and compute a shadow deviation distance $\delta_{s}$ on shadow model based on the same known features, following the same reconstruction procedure. In summary, both $\delta$ and $\delta_{s}$ are derived using identical feature values, but from the target and shadow models respectively. The shadow deviation $\delta_{s}$ serves as a reference for how typical the deviation would be if the known features is not a member of the training set. Then, we define a membership inference rule based on the relative deviation ${\delta_{s}}/{\delta}$. Let MRAD denote this relative deviation and the inference rule can be formulated as:
@@
Threshold Selection: Previous work has shown that MIA is only meaningful when the false positive rate is sufficiently low[carlini2022membership]. Therefore, we adopt a threshold selection strategy similar to that proposed by Ye et al.[ye2022enhanced]. Specifically, the adversary can sample a batch of samples $\mathcal{D}_{non}$ from the overall non-member data space $\pi^{\prime}$. Given a predefined acceptable false positive rate $\alpha$, the adversary then selects a threshold that satisfies
@@
IV-C Black-box Attack
The white-box results establish the core feasibility of our approach, but this setting rarely reflects practical deployment environments. To bridge this gap, we develop a black-box variant of our attack. The key idea is to approximate the required gradients using zeroth-order gradient estimation[wibisono2012finite], allowing the attacker to operate without direct access to the model’s internal parameters or intermediate computations.
Zeroth-order gradient estimation approximates the gradient via finite differences: small perturbations are applied to the input along multiple random directions $\mu_{i}$, and the corresponding output variations are used to construct the gradient estimate. Let $m$ denote the number of random directions, $\epsilon$ the perturbation magnitude. $\mathcal{L}$, as the loss function, serves as the objective function for gradient estimation. Then, the estimated gradient at a $p$-dimensional data point $x$ whose loss under the loss function $L$ can be formulated as
$\nabla\mathcal{L}=\frac{1}{m}\sum_{i=1}^{m}p\frac{f(x+\epsilon\mu_{i})-f(x)}{\epsilon}\mu_{i}.$ (4)
Since zeroth-order gradient estimation requires drawing multiple samples in the neighborhood of the target point $x$, the black-box attack issues $m+1$ times more queries than its white-box counterpart—$m$ for perturbed inputs and one for the original sample. A very large $m$ not only increases query overhead but also makes the query pattern more likely to be interpreted as unusual or suspicious behavior by the target system, whereas an overly small $m$ leads to a poor approximation of the underlying gradient. Consequently, $m$ acts as a key trade-off variable that balances query efficiency and stealthiness against the fidelity of the estimated gradient and, ultimately, overall attack performance. We will test the black-box setting at section V-F.
@@ V Evaluation
In this section, we evaluate the attack performance of the proposed algorithm. We begin with a quantitative assessment under the white-box setting using four widely adopted datasets. These experiments examine the capability of our attack across different known-feature percentage and include a series of in-depth analyses, such as parameter sensitivity experiments, ablation study, and evaluations of feature-importance effects. We then test the black-box variant of our algorithm on the same four datasets to assess its effectiveness under restricted access conditions. Finally, we demonstrate how our attack operates in real-world scenarios to highlight its practical applicability.
@@ V-A Experiment setup
Dataset & Model: We evaluate our attack with three widely used image datasets and one tabular dataset. For all image datasets, the target model is Resnet50, while for the tabular dataset, we employ a multilayer perceptron (MLP). We train each model for 150 epochs, with a learning rate of 0.01 for image datasets and 0.1 for tabular datasets. We provide a detailed overview of each dataset as below.
  • CIFAR-10[krizhevsky2009learning]: CIFAR-10 is a widely used benchmark dataset for image classification. It contains 50,000 training images and 10,000 test images, each being a 32×32 RGB image.
  • Fashion-MNIST[xiao2017fashion]: Fashion-MNIST is a drop-in replacement for the original MNIST dataset, featuring 60,000 training images and 10,000 test images. Each image is a 28×28 grayscale representation of an article of clothing.
  • STL-10[coates2011analysis]: STL-10 contains 13,000 labeled images and 100,000 unlabeled images, making it suitable for supervised, unsupervised, and self-supervised learning. In our experiments, we use only the labeled subset. Each image is a 96×96 RGB image.
  • Epsilon[pascal_large_scale]: Epsilon was derived from the PASCAL 2008 challenge, contains simulated physical data for a binary classification task. Each sample consists of 2,000 normalized numerical features.
@@
Anomaly Detection Methods: In addition to our proposed method, we integrate three representative anomaly detection techniques: CADE [yang2021cade], MSAD [reiss2023mean], and NCI [liu2025detecting]. These comparisons demonstrate the compatibility of our framework with various existing techniques and allow for a comprehensive evaluation across different methodologies. The three baseline methods are summarized as follows:
@@
  • NCI: NCI is a post-hoc OOD detection method based on the phenomenon of neural collapse. It observes that in-distribution (ID) samples tend to align with their class prototype vectors in the penultimate layer, whereas OOD samples do not. Additionally, ID features tend to cluster near the origin. We adopt the publicly available implementation from github[zhang2023openood].
V-B Performance Evaluation
We begin by evaluating the attack performance under different proportions of known features. We vary the known feature percentage from 10% to 90% in increments of 10% and conduct experiments at each level. In addition, we include a leave-one-out setting where only one feature is missing. This setting is not intended to model a realistic attacker, but rather serves as an extreme and informative limit case of partial-feature knowledge. Because the gradient magnitude is small, we set a larger step size of $\eta=1000$ and perform only one optimization step (the effect of hyperparameters will be further discussed in Section V-C). The results are shown in Figures 3 and 4, where the x-axis indicates the percentage of known features. The leave-one-out setting is marked as ”loo” on the x-axis.
The evaluation results demonstrate a clear pattern. The attack becomes increasingly effective as more features are known to the adversary. On CIFAR-10 and STL-10, the AUC remains above 0.7 even when half of the features are missing, and the attack still achieves a strong TPR. The Epsilon dataset follows a similar trajectory. Although its absolute performance is slightly lower than that of the image datasets, the attack remains consistently effective and well within a practically usable range. When the known-feature ratio reaches 90%, attack performance peaks on most datasets, with all image datasets achieving an AUC of at least 0.8. This upward trend of attack performance is expected, since additional known features provide tighter constraints for gradient-based reconstruction, enabling the attacker to produce reconstructions that are more consistent with the true data distribution.
However, an important reversal occurs when the known-feature ratio becomes exceedingly high. As more features are revealed, regardless of how poorly the unknown features are reconstructed, the samples increasingly align with the original data distribution. Even when the known features are from a non-member, the reconstructed samples may still evade detection by the anomaly detection algorithm in the second stage. Consequently, the overall attack performance degrades. When the known-feature ratio reaches 100%, Stage I becomes redundant and the method effectively reduces to a conventional membership inference attack. In that extreme case, standard MIA techniques are typically the more appropriate choice.
@@ V-B Performance Evaluation
V-C Hyperparameter Sensitivity
Here we further analyze the influence of two key hyperparameters with a known percentage of 80%. For each experimental setting, we adopt NCI as the anomaly detection method and repeat the attack ten times and report the average performance together with its range. All values are rounded to two decimal places, which may cause very small numbers to appear as 0.00.
Figure 5, Figure 6, Figuure 7 and Figure8 report the results of applying our attack to four datasets. Based on the experimental results, we identify two parameter settings that can lead to strong attack performance.
The first approach is to use a large step size $\eta$ together with a small number of optimization iterations $N_{itr}$. As shown in the heatmaps, when $N_{itr}$ is fixed, increasing $\eta$ generally improves the attack performance within the range we evaluated. This is because, for a sample that follows the normal data distribution, the loss produced by a well-trained target model remains small even when some features are missing. A larger step size is therefore necessary to amplify the gradient signal and make the optimization updates effective. In addition, when $\eta$ is sufficiently large, only a few optimization iterations are needed to achieve strong results, while excessively increasing $N_{itr}$ may even reduce the attack effectiveness.
(a) AUC
(b) TPR@0.1FPR
Figure 5: Attack performance on the CIFAR-10 dataset under different parameter settings. Darker blue indicates higher attack performance. The number in each cell denotes half of the range (i.e., half the max–min difference) across ten runs.
(a) AUC
(b) TPR@0.1FPR
Figure 6: Attack performance on the STL-10 dataset under different parameter settings. Darker blue indicates higher attack performance. The number in each cell denotes half of the range (i.e., half the max–min difference) across ten runs.
(a) AUC
(b) TPR@0.1FPR
Figure 7: Attack performance on the Fashion-MNIST dataset under different parameter settings. Darker blue indicates higher attack performance. The number in each cell denotes half of the range (i.e., half the max–min difference) across ten runs.
(a) AUC
(b) TPR@0.1FPR
Figure 8: Attack performance on the Epsilon dataset under different parameter settings. Darker blue indicates higher attack performance. The number in each cell denotes half of the range (i.e., half the max–min difference) across ten runs.
The second approach is to choose a moderate step size $\eta$ and perform multiple rounds of optimization (e.g., $N_{itr}=20,\eta=10$ on CIFAR-10 or $N_{itr}=20,\eta=100$ on STL-10). However, this strategy does not outperform the first approach and increases the overall computational cost—especially in black-box settings, where each optimization step requires $(m+1)$ model queries. For this reason, we recommend using a relatively large $\eta$ with only a single optimization iteration, which provides a more efficient and stable configuration.
V-D Ablation Study